I need to solve :
1/2 + x/3 = 2x/3 - 1/4
Thanks for your help

Answer by guapa(62)

1/2 +x/3 = (2x)/3 - 1/4
First you need to find the LCD. In doing it you get rid of all fractions in the equation. In this case the LCD is 12.
(1*6)/(2*6) + (x*4)/(3*4) = (2x*4)/(3*4)- (1*3)/(4*3)
6 + 4x = 8x -3
4x = 9
x = 9/4